OpenStudy (pulsified333):

Assume that the class consists of 55 percent freshmen, 5 percent sophomores, 25 percent juniors, and 15 percent seniors. Assume further that 55 percent of the freshmen, 40 percent of the sophomores, 20 percent of the juniors, and 20 percent of the seniors plan to go to medical school. One student is selected at random from the class. (1) What is the probability that the student plans to go to medical school? .4125 (2) If the student plans to go to medical school, what is the probability that he is a sophomore? .02/.4125

OpenStudy (pulsified333):

Tell me why the answer I got are wrong please

OpenStudy (pulsified333):

@dan815

OpenStudy (badhi):

can you tell us how you tried to solve this problem?

OpenStudy (pulsified333):

using a tree but it obviously did not work

OpenStudy (badhi):

first try to write the given information in a notatio.. freshman = f junior = j sophomore = so senior = s the following are given,\[P(f), P(s), P(so), P(j)\] And also if selecting medicine is m \[P(m|j) = 0.2, P(m|s) = 0.2\] The question asks for probability of the selected student hopes to do medicine -> P(m) use the conditional probability equations to obtain that

OpenStudy (pulsified333):

I did that

OpenStudy (pulsified333):

I did (.55*.55)+(.05*.4)+(.25*.2)+(.2*.2)

OpenStudy (badhi):

Is it wrong?

OpenStudy (pulsified333):

wait i think i see my mistake

OpenStudy (pulsified333):

wait I have no clue what I did

OpenStudy (badhi):

\[P(m) = P(m\cap s) + P(m \cap j) +\cdots\\ = P(m|s)P(s) + P(m|j)P(j)+ \cdots\] so i think its same as what youve done the answer should be correct.. does it give a correct answer?

OpenStudy (pulsified333):

yeah when I get the correct answer but it isn't

OpenStudy (badhi):

what does it state as the correct answer?

OpenStudy (pulsified333):

It doesn't. It will only say if its correct when I get the correct answer

OpenStudy (badhi):

I think your answer is correct :(

OpenStudy (kropot72):

@Pulsified333 "I did (.55*.55)+(.05*.4)+(.25*.2)+(.2*.2)" The last term in brackets should be .......+(.15 * .2). With this correction the sum of the terms is 0.4025 which is the probability that the student plans to go to medical school.

OpenStudy (pulsified333):

why (.15*.2)

OpenStudy (pulsified333):

@dan815

OpenStudy (pulsified333):

its the correct answer but why?

OpenStudy (kropot72):

The question states that 15% of the class is seniors, and 20% of the seniors plan to go to medical school. The intersection of P(senior) and P(med.school|senior) = .15 * .2.

OpenStudy (pulsified333):

oh! that makes sense now :D

OpenStudy (kropot72):

This is the same way that you have correctly calculated the values of the other three intersections.

OpenStudy (pulsified333):

oh okay

OpenStudy (kropot72):

You're welcome :)

OpenStudy (pulsified333):

thank you